To get to work, you drive for 20 minutes at 30 mph, and walk 10 more minutes at 3 mph. What was your average speed?

a) 16.5 mph
b) 20 mph
c) 21 mph
d) 23 mph

Answer by ewatrrr(24785)

 
Hi,
Note D = r*t 20min = 1/3 hr and 10min = 1/6 hr
(1/3hr)30mph + (1/6hr)3mph = 10.5 mi, the total distance traveled.
total distance / total time = average speed = 10.5/(1/3 + 1/6) = 10.5/(2/6 + 1/6)
10.5mi/.5hr = 21mph